Determine the sum of the geometric series 27+18+12+8+...
Answer by fractalier(2101)

Here a1 = 27 and r = 2/3...the formula is
S = a1 / 1-r
S = 27 / (1-2/3)
S = 27 / (1/3)
S = 81